# Keep it Light! Simplifying Image Clustering via Text-Free Adapters

### v1

@@ 4.3 Visualization
Figure 3: The visualization of clustering performance for SCP-CLIP with ViT-B/32 backbone. (Left): An example of an image-to-image search on STL-10, showing clusters produced by CLIP (Top) and SCP (Bottom); (Right): Visualization of clustering performance. SCP-CLIP effectively enhances CLIP’s clustering performance through a cluster head.
To provide an intuitive understanding of the clustering results, we visualize the clustering performance obtained from SCP in Fig. 3, along with learned representations on the CIFAR-10 dataset in Fig. 4. The image logits, representing SCP outputs before the final $\operatorname{softmax}$ function, are used for visualization, with t-SNE applied to reduce the feature dimensions. As shown, compared to embeddings directly learned by pre-trained models, SCP effectively forms well-separated clusters, leading to a higher ACC score. Without aligned text, SCP successfully extracts image features by incorporating a clustering head, resulting in superior within-cluster compactness and between-cluster separability. The visualization supports our quantitative results and highlights SCP’s effectiveness in learning discriminative features suitable for lightweight clustering tasks. Furthermore, Fig. 2 demonstrates that SCP requires only a few steps to establish a clustering structure, effectively separating certain image embeddings within 100 steps. For instance, from step 10 to step 50, SCP learns to push the light-green cluster away from the blue cluster.
@@
Figure 4: Visualization of representations learned by different methods on the CIFAR-10 training set, along with the corresponding clustering accuracy (ACC). (a) Image embeddings directly from the DINO image encoder, with clusters obtained by K-means. (b) Image logits from SCP-DINO. (c) Image embeddings directly from the CLIP image encoder, again clustered by K-means. (d) Image logits from SCP-CLIP.
@@ 4.4 Ablation Studies
To assess the effectiveness of the three loss terms, $L_{\text{e}}$, $L_{\text{con}}$, and $H(Y)$, we evaluate the performance of SCP using different combinations of these losses, as presented in Table 4. The results reveal several important insights: i) $H(Y)$ plays a crucial role in preventing cluster collapse. Without $H(Y)$, SCP tends to assign most images to only a few clusters, resulting in the model collapse on CIFAR-10 and CIFAR-20. ii) $L_{\text{con}}$ provides a slight boost in performance. The reason is that the cluster assignments would be less confident when the cluster number increases. iii) The combination of all losses effectively learns image information, leading to the best clustering performance in this table.
Table 4: Clustering results for different losses on CIFAR-10 and CIFAR-20. A ’-’ indicates model collapse.
@@ 4.5 Parameter Analyses
To show how the scale of $H(Y)$ influences the performance of SCP, we evaluate it under various choices of $\alpha$ on training sets of CIFAR-10, CIFAR-20, CIFAR-100 and ImageNet-Dogs under 5 random seeds. The average results and the corresponding standard deviations are presented in Fig. 5.
Figure 5: Comparison of different loss weights $\alpha$, the solid line is the mean ARI across five runs, and the shaded region plots the standard deviation across our runs.
We observe that setting the loss weight $\alpha$ too low (e.g., below 0.5) makes it difficult for our method to converge. For CIFAR-10, the performance remains relatively stable and is not sensitive to large $\alpha$ values. In contrast, for ImageNet-dogs and CIFAR-20, choosing $\alpha=2$ leads to better results, suggesting it provides an optimal balance between regularization and scale. For CIFAR-100, which involves a larger number of clusters, a higher $\alpha$ value is more suitable. Based on these findings, we select $\alpha=3$ for CIFAR-100, $\alpha=2$ for CIFAR-20 and ImageNet-dogs, and $\alpha=1$ for all other datasets. It is also consistent with the prior knowledge of the number of clusters in each dataset.
@@ 5 Conclusion
In this work, we set out to determine whether a simple DC pipeline could achieve the performance of other recent clustering frameworks. Specifically, we aimed to achieve this without relying on text, as we argue that doing so would make image-based clustering much more practical for downstream applications. This is because image-pair datasets are highly unlikely to be readily available to practitioners for most real-world use cases. Moreover, we aimed to develop a method that is simple to train without the need for additional feature layers, support sets, teacher-student networks, or exponential moving averages.
To meet this desiderata we proposed SCP, a novel, simple, and lightweight clustering method that achieves SOTA and near-SOTA performance in several datasets such as STL-10, CIFAR-10, CIFAR-20, ImageNet-10, and ImageNet-Dogs. In particular, SCP leverages pre-trained CLIP and DINO backbones, and we anticipate these could be updated in the future with better models, or ensembles.
Limitations
Despite these promising results, several challenges remain. Firstly, our pipeline relies on data augmentations, which can be time-consuming when scaling to larger datasets such as ImageNet-1K. Also, our approach requires prior knowledge of the number of clusters, which can be difficult to determine in practical applications unless the practitioner poses strong domain knowledge. Lastly, our frameworks performance is bounded by the quality of the pre-trained model used. However, we anticipate these to continue to improve in the future.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Additional Experiment Details
A.1 Image Search Pipeline
A.1.1 CLIP for Image search
Figure 6: CLIP Image-to-Image Search baseline depicted in Fig. 3.
Figure 6 illustrates the CLIP baseline for image-to-image search. The image repository comprises the STL-10 test split, with a target image randomly selected for the search. The search process ranks images based on Euclidean distances, displaying the top 20 most similar results. The results indicate that CLIP effectively learns discriminative features, although two mismatched images are still present in the top-20 neighborhood.
A.1.2 SCP for Image search
Figure 7 illustrates the SCP-CLIP approach for image-to-image search. SCP-CLIP successfully retrieves images from the same cluster within the top-20 neighborhood.
Figure 7: SCP-CLIP Image-to-Image Search Pipeline depicted in Fig. 3, utilizing the normalized outputs of the trained cluster head $g(.)$ without applying softmax.
A.2 Visualization
We create a t-SNE visualization image cloud of STL-10 test set to support SCP effectiveness further. It was observed that both SCP-DINO and SCP-CLIP successfully distinguished most planes, ships, horses, and deer, even when they share a blue background or similar body shape.
Figure 8: The t-SNE representations learned by SCP-DINO on STL-10. $ACC:98.4\%$
Figure 9: The t-SNE representations learned by SCP-CLIP on STL-10. $ACC:97.9\%$
Appendix B Proofs
Proof of Proposition 1.1.
By the Doob-Dyknin lemma, see e.g. [49, Lemma 1.14], since $Z$ is $\sigma(X)$-measurable if and only if there exists some Borel measurable function $g:(\mathbb{R}^{d},\mathcal{B}(\mathbb{R}^{d}))\to(\mathbb{R}^{D},\mathcal{B}(\mathbb{R}^{D}))$ such that
$Z=g(X).$ (7)
Now, define $F:(\mathbb{R}^{d},\mathcal{B}(\mathbb{R}^{d}))\to(\mathbb{R}^{n},\mathcal{B}(\mathbb{R}^{n}))$ by composition, as sending every $x\in\mathbb{R}^{d}$ to
[…]
Then, together (7), the definition of $Y$ in (1), and that of $F$ in (8) yield
$Y=f(X,Z)=f(X,g(X))=F(X)$ (9)
which establishes our claim. ∎
Proof of Theorem 1.2.
[…]
$\mu(\mathcal{K}_{\varepsilon})\geq 1-\varepsilon$ (10)
and for which the restriction of the Borel measurable map $F$ to $\mathcal{K}_{\varepsilon}$ is continuous.
Now, fix a continuous activation function $\rho:\mathbb{R}\to\mathbb{R}$ with at least one point of continuous non-zero differentiability; i.e. satisfying [51, Assumption 1] (due to [52]). As shown in [51, Example 13], the softmax function satisfies [51, Assumption 8]; whence the general case of the non-Euclidean universal approximation theorem [51, Theorem 37 (ii)] applies; from which we deduce the existence of an MLP $\hat{F}:\mathbb{R}^{d}\to\mathbb{R}^{C}$ with $\rho$ activation function satisfying the uniform approximation guarantee
$\sup_{x\in\mathcal{K}_{\varepsilon}}\,|F(x)-\operatorname{softmax}\circ\hat{F}(x)|<\varepsilon.$ (11)
Define the failure set
[…]
Consequentially, the following computations are well founded: Combining (10) and (11) we deduce that
$\mathbb{P}\biggl{(}|F(X)-\operatorname{softmax}\circ\hat{F}(X)|<\varepsilon\biggr{)}$ $=\mu\biggl{(}|F(x)-\operatorname{softmax}\circ\hat{F}(x)|<\varepsilon\biggr{)}$ (12)
$=\mu(\mathbb{R}^{d}\setminus B)$
$\geq\mu(\mathcal{K}_{\varepsilon})$
$\geq 1-\varepsilon.$ (13)
Finally, by Proposition 1.1 and (1), we know that
$F(X)=f(X,Z).$ (14)
Whence, incorporating the identity in (14) into the left-hand side of the chain of inequalities in (12)-(13) implies that
$\mathbb{P}\biggl{(}|f(X,Z)-\operatorname{softmax}\circ\hat{F}(X)|<\varepsilon\biggr{)}=\mathbb{P}\biggl{(}|F(X)-\operatorname{softmax}\circ\hat{F}(X)|<\varepsilon\biggr{)}\geq 1-\varepsilon$ (15)
which concludes our proof. ∎

### v2

@@ 4.3 Visualization
Figure 2: The visualization of clustering performance for SCP-CLIP with ViT-B/32 backbone. (Left): An example of an image-to-image search on STL-10, showing clusters produced by CLIP (Top) and SCP (Bottom); (Right): Visualization of clustering performance. SCP-CLIP effectively enhances original CLIP’s clustering performance.
To provide an intuitive understanding of the clustering results, we visualize the clustering performance obtained from SCP in Fig. 2, along with learned representations on the CIFAR-10 dataset in Fig. 3. The image logits, representing SCP outputs before the final $\operatorname{softmax}$ function, are used for visualization, with t-SNE applied to reduce the feature dimensions. As shown, compared to embeddings directly learned by pre-trained models, SCP effectively forms well-separated clusters, leading to a higher ACC score. Without aligned text, SCP successfully extracts image features by incorporating a clustering head, resulting in superior within-cluster compactness and between-cluster separability. The visualization supports our quantitative results and highlights SCP’s effectiveness in learning discriminative features suitable for clustering tasks. Furthermore, Fig. 4 demonstrates that SCP requires only a few steps to establish a clustering structure, effectively separating certain image embeddings within 100 steps. For instance, from step 10 to step 50, SCP learns to push the light-green cluster away from the blue cluster.
@@
Figure 3: Visualization of representations learned by different methods on the CIFAR-10 training set, along with the corresponding clustering accuracy (ACC). (a) DINO + K-means. (b) CLIP + K-means. (c) SCP-DINO logits. (d) SCP-CLIP logits.
(a) Step 1, ARI = 0.045
(b) Step 10, ARI = 0.24
(c) Step 50, ARI = 0.57
(d) Step 100, ARI = 0.59
Figure 4: Visualization of learned representations at different training steps on ImageNet-Dogs. Each panel shows the t-SNE embeddings of the DINO-based features, where (a) depicts the raw encoder outputs clustered by K-means, and (b)–(d) show the logits learned by SCP-DINO across successive training stages (77 steps per epoch).
@@ 4.4 Ablation Studies
TABLE VII: Clustering results for different losses on CIFAR-10 and CIFAR-20. A “–” indicates model collapse.
@@
To assess the effectiveness of the three loss terms, $L_{\text{e}}$, $L_{\text{con}}$, and $H(Y)$, we evaluate the performance of SCP using different combinations of these losses, as presented in Table VII. The results reveal several important insights: i) $H(Y)$ plays a crucial role in preventing cluster collapse. Without $H(Y)$, SCP tends to assign most images to only a few clusters, resulting in the model collapse on CIFAR-10 and CIFAR-20. ii) $L_{\text{con}}$ provides a slight boost in performance. The reason is that the cluster assignments would be less confident when the cluster number increases. iii) The combination of all losses effectively learns image information, leading to the best clustering performance in this table.
@@ 4.5 Parameter Analyses
To show how the scale of $H(Y)$ influences the performance of SCP, we evaluate it under various choices of $\alpha$ on training sets of CIFAR-10, CIFAR-20, CIFAR-100 and ImageNet-Dogs under 5 random seeds. The average results and the corresponding standard deviations are presented in Fig. 5. We observe that setting the loss weight $\alpha$ too low (e.g., below 0.5) makes it difficult for our method to converge. For CIFAR-10, the performance remains relatively stable and is not sensitive to large $\alpha$ values. In contrast, for ImageNet-dogs and CIFAR-20, choosing $\alpha=2$ leads to better results, suggesting it provides an optimal balance between regularization and scale. For CIFAR-100, which involves a larger number of clusters, a higher $\alpha$ value is more suitable. Based on these findings, we select $\alpha=3$ for CIFAR-100, $\alpha=2$ for CIFAR-20 and ImageNet-dogs, and $\alpha=1$ for all other datasets. It is also consistent with the prior knowledge of the number of clusters in each dataset.
Figure 5: Comparison of different loss weights $\alpha$. The solid line shows the mean ARI across five runs, and the shaded region indicates the standard deviation.
Limitations
Despite these promising results, several challenges remain. Firstly, our pipeline relies on data augmentations, which can be time-consuming when scaling to larger datasets such as ImageNet-1k. Also, our approach requires prior knowledge of the number of clusters, which can be difficult to determine in practical applications unless the practitioner possesses strong domain knowledge. Lastly, our framework’s performance is bound by the quality of the pre-trained model used. However, we anticipate these to continue to improve in the future.
@@ 5 Conclusion
In this work, we investigate whether SOTA performance can be achieved using a simple deep clustering method. This investigation is motivated by the observation that many current methods rely on text or extra frameworks, making training complicated and widespread adoption challenging. We first provide a theoretical justification, arguing that strong performance can be achieved using informative visual representations alone. Based on this, we propose SCP, a simple and text-free adapter that requires no additional components during training, aiming for effective and accessible clustering. By leveraging pre-trained models, SCP-CLIP already demonstrates competitive performance on CIFAR-10, CIFAR-20, and STL-10. To further support our claim, we introduce SCP-DINO and SCP-MIX, which leverage stronger visual representations and achieve SOTA results across multiple benchmarks. Our findings demonstrate that SCP provides a practical and effective alternative for image clustering, featuring a simple design, easy deployment, broad accessibility, and competitive performance.
Acknowledgments
The authors gratefully acknowledge extensive computational support from the Digital Research Alliance of Canada. This work was supported by a Dorothy Killam Fellowship, the Canada Research Chairs program, and an NSERC Discovery Grant.